• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 23:54:32 on June16,2002

    Entry number 86201

    keyword=Shift Summary

    17:15 MCC called to inform us that they want to increase raster at a higher rate.
    16:30 ROC14 hanged.
    19:50 ROC14 hanged again.
    20:00 ROC14 hanged again.
    21:00 When switching from 15 cm LD2 to 4 cm LH2, we ran into slight
    problem. The fan speed readouts between screen and software
    don't agree. Called J. P. and he told us to use screen readout.

    Beam has been very stable at 100 uA through the entire shift! It has been a very
    busy shift; there are many spectrometer momentum and angle changes, plus
    target and collimator changes. In addition, runs are short in duration. During
    this shift, we finish kinematics Q2_dfN and Q2_f00, and are half way through
    Q2_Acc. Shift has been quite un-eventfull except a few ROC14 problem.